# Tate modules as condensed modules

### v1

@@ 3. $\infty$-categorical embedding in the bounded case
Although we will not use it, we note that the proof of lemma 3.6 does not make use of the boundedness assumption ($\star$3.4).
We now establish that the comparison morphism $\gamma$ is surjective on connected components, as stated more precisely in the following result.
Lemma 3.8.
Let $f\colon\lim V\to W$ be a morphism in $\mathscr{D}(\mathsf{Cond}_{R})$, with $W$ a discrete module. Assume that the boundedness hypothesis (italic-$\star$3.4). Then $f$ factors through a projection $p_{i}\colon\lim V\to V_{i}$, for some $i\in I$.
Before proving the lemma, we introduce some convenient constructions and notations.
Notation 3.9.
For $n\in\mathbb{Z}$, we consider the functor
$\Omega^{\infty+n}:=\operatorname{Map}_{\mathscr{D}(R)}(\Sigma^{n}R,-)\colon\mathscr{D}(R)\longrightarrow\mathscr{S}$
which is right adjoint to the free functor $\Sigma^{\infty}(-)\otimes\Sigma^{n}R$. Since it commutes with limits, it induces a functor $\mathscr{D}(\mathsf{Cond}_{R})\to\mathsf{Cond}(\mathscr{S})$ sending a condensed $R$-module $A$ to the condensed pointed $\infty$-groupoid $S\mapsto\Omega^{\infty+n}(A(S))$; we still denote this new functor by $\Omega^{\infty+n}$.
Notation 3.10.
Given integers $a$ and $b$, we consider the functor
$\mathrm{U}:=\prod_{n\in[a,b]}\Omega^{\infty+n}\colon\mathscr{D}(\mathsf{Cond}_{R})\longrightarrow\mathsf{Cond}(\mathscr{S})$
leaving the dependence on $a$ and $b$ implicit in the notations.
Lemma 3.11.
If $W\in\mathsf{Cond}(\mathscr{S})$ is a discrete condensed $\infty$-groupoid, then for every profinite sets $S$ we have an equivalence
$W(S)\simeq\operatorname*{colim}_{\text{covers }\mathcal{U}\text{ of }S}W(*)^{|\mathcal{U}|}.$ (3.4)
Proof.
The constant sheaf $W$ is given by the sheafification $L\mathrm{const}_{W(*)}$ of the constant presheaf with value $W(*)\in\mathscr{C}$. Following [Lur09, Proposition 6.2.2.7.] and its proof, the sheafification functor can be constructed as a transfinite composition of the functor $(-)^{\dagger}\colon\mathscr{P}(\mathsf{ProFin})\to\mathscr{P}(\mathsf{ProFin})$ given on objects by the formula
$\mathcal{F}^{\dagger}(S)\simeq\operatorname*{colim}_{\mathcal{U}}\lim_{S^{\prime}\in\mathcal{U}}\mathcal{F}(S^{\prime})$ (3.5)
where the colimit runs over all covers $\mathcal{U}$ of $S$. More precisely, there exists a regular cardinal $\kappa$ and a transfinite sequence $\alpha\mapsto T_{\alpha}$ of endofunctors of $\mathscr{P}(\mathsf{ProFin})$, indexed by the linear poset of ordinals $\alpha\leq\kappa$, such that $L\simeq T_{\kappa}$ and all maps $T_{\alpha}\to T_{\alpha^{\prime}}$ for $\alpha<\alpha^{\prime}$ are local equivalences. The sequence can be defined using transfinite induction by $T_{0}=\mathrm{id}$, $T_{\alpha+1}=T_{\alpha}^{\dagger}$ and for any limit ordinal $\beta$, $T_{\beta}=\operatorname*{colim}_{\alpha<\beta}T_{\alpha}$. Since $(-)^{\dagger}\simeq T_{1}\to T_{\kappa}\simeq L$ is a local equivalence and the right hand side of equation (3.4) is precisely $(\mathrm{const}_{W(*)})^{\dagger}$, it suffices to show that the latter presheaf is actually a sheaf.
Let $S_{1}$ and $S_{2}$ be two extremally disconnected sets. We want to prove that the natural map
$\mathrm{const}_{W(*)}^{\dagger}(S_{1}\amalg S_{2})\longrightarrow\mathrm{const}_{W(*)}^{\dagger}(S_{1})\times\mathrm{const}_{W(*)}^{\dagger}(S_{2})$
is an equivalence. Unraveling the definitions, the above morphism corresponds to the canonical map
$\operatorname*{colim}_{\text{covers }\mathcal{U}\text{ of }S_{1}\amalg S_{2}}W(*)^{|\mathcal{U}|}\longrightarrow\mathopen{}\mathclose{{}\left(\operatorname*{colim}_{\text{covers }\mathcal{U}_{1}\text{ of }S_{1}}W(*)^{|\mathcal{U}_{1}|}}\right)\times\mathopen{}\mathclose{{}\left(\operatorname*{colim}_{\text{covers }\mathcal{U}_{2}\text{ of }S_{2}}W(*)^{|\mathcal{U}_{2}|}}\right).$
But the latter is easily seen to be an equivalence since filtered colimits commute with products and covers of $S_{1}\amalg S_{2}$ are exactly pairs of covers of $S_{1}$ and $S_{2}$. ∎
Remark 3.12 (Topological inspiration for the proof of lemma 3.8).
Before showing the lemma, let us explain a proof of the topological analog of this result, namely the fact that any map of topological $R$-modules $f\colon\lim_{i}V_{i}\to W$, from the limit of a countable sequential pro-object $V$ to a discrete module $W$, factors through one of the $V_{i}$’s.
Assume towards a contradiction that for every $i\in I$, we can pick a vector $v_{i}\in\ker(p_{i})$ that does not belong to $\ker(f)$. The sequence $(v_{i})_{i\in I}$ tends to 00 as $i\to\infty$, therefore so does its image $f(v_{i})$ in $W$. Since $W$ is discrete, this sequence must be eventually null; in particular, $f(v_{k})=0$ for some $k\in\mathbb{N}$, which then contradicts $v_{k}\notin\ker(f)$. Hence $f$ must factor through one of the $V_{i}$’s.
The actual proof in $\mathscr{D}(\mathsf{Cond}_{R})$ given below is an adaptation of this topological argument.
Proof of lemma 3.8.
Assume towards a contradiction that for every $i\in I$, the morphism $f\circ\iota_{i}$ is nonzero in $\pi_{0}\operatorname{Map}(\mathrm{fib}(p_{i}),W)$. Choose a regular cardinal $\kappa$ large enough so that all the $V_{i}$ are $\kappa$-condensed objects. The set of $\Sigma^{n}R[S]$ for varying $\kappa$-small profinite sets $S$ and $n\in\mathbb{Z}$ generate $\mathscr{D}(\mathsf{Cond}^{\kappa}_{R})$ under small colimits, therefore for every $i\in I$ we may write $\mathrm{fib}(p_{i})$ as a small colimit of such generators; since $f\circ\iota_{i}\not\simeq 0$, we can pick some $\kappa$-small non-empty profinite set $S_{i}$, some integer $n_{i}$ and a morphism $v_{i}\colon\Sigma^{n_{i}}R[S_{i}]\to\mathrm{fib}(p_{i})$ such that $f\circ\iota_{i}\circ v_{i}\not\simeq 0$.
We claim that the profinite sets $S_{i}$ can be chosen independently of $i$. To see this, consider their product $S=\prod_{i\in I}S_{i}$. Note that each projection $S\to S_{i}$ admit a section so the induced map $q_{i}\colon\Sigma^{n_{i}}R[S]\to\Sigma^{n_{i}}R[S_{i}]$ is an epimorphism. It follows that the maps $\widetilde{v}_{i}=v_{i}\circ q_{i}$ still satisfy $f\circ\iota_{i}\circ\widetilde{v}_{i}\not\simeq 0$.
By the boundedness assumption ($\star$3.4), we can pick $a,b\in\mathbb{Z}$ such that either $V$ or $W$ is of tor-dimension in $[-b,-a]$, so that $n_{i}\in[a,b]$ for every $i\in I$ (recall that we use cohomological grading conventions). We will write $v^{\prime}_{i}$ for the composite
$v^{\prime}_{i}\colon\bigoplus_{n\in[a,b]}\Sigma^{n}R[S]\xrightarrow{\mathrm{proj}_{n_{i}}}\Sigma^{n_{i}}R[S]\xrightarrow{\widetilde{v}_{i}}\mathrm{fib}(p_{i})\xrightarrow{\iota_{i}}\lim V,$
which we can view equivalently as an element $v^{\prime}_{i}\in\mathrm{U}\lim V(S):=\prod_{n\in[a,b]}\Omega^{\infty+n}\lim V(S)$. Here we made use of the functor $\mathrm{U}$ defined in (3.10), implicitly depending on the interval $[a,b]$. Note that this element $v^{\prime}_{i}$ satisfies
$f(v_{i}^{\prime})\not\simeq 0\quad\text{ in }\quad\mathrm{U}W(S).$ (3.6)
The next step is to assemble the different sections $v_{i}^{\prime}$ into a single "converging sequence", in the sense of an element $v\in\mathrm{U}\lim V(S\times\mathbb{N}^{\infty})$, where $\mathbb{N}^{\infty}$ denotes the profinite set obtained as the one-point compactification of the discrete set $\mathbb{N}$. Since $\mathrm{U}$ commutes with limits, we obtain an equivalence $\mathrm{U}\lim_{j}V_{j}\simeq\lim_{j}\mathrm{U}V_{j}$ of condensed $\infty$-groupoids. Therefore, to construct the desired sequence $v$ it suffices to produce a compatible sequence of elements $v(j)$ in $\mathrm{U}V_{j}(S\times\mathbb{N}^{\infty})$, for varying $j\in I$. We proceed as follows. By construction, we have $p_{j}v_{i}^{\prime}\simeq 0$ for $i\geq j$ so that for a fixed $j\in I$, the sequence $(p_{j}v_{i}^{\prime})_{i\in I}$ in $V_{j}(S)$ is eventually null. Using the cover
$(S\times\{0\})\amalg\dots\amalg(S\times\{j-1\})\amalg(S\times\llbracket j,\infty\rrbracket)\longrightarrow S\times\mathbb{N}^{\mathrm{\infty}}$
in $\mathsf{ProFin}$, the local sections $(p_{j}v_{0}^{\prime},\dots,p_{j}v_{j-1}^{\prime},0)$ glue into a section $v(j)\in\mathrm{U}V_{j}(S\times\mathbb{N}^{\infty})$. Observe that for any $j\geq j^{\prime}$ these sections come with equivalences $p_{jj^{\prime}}v(j^{\prime})\simeq v(j)$, where $p_{jj^{\prime}}$ denotes the natural map $V_{j}\to V_{j^{\prime}}$. More generally, the $v(j)$’s are compatible in the following sense: they are part of a natural transformation $\mathrm{const}_{S\times\mathbb{N}^{\infty}}\to\mathrm{U}V$ of functors $I\to\mathsf{Cond}(\mathscr{S})$ from the constant diagram with value the representable sheaf $S\times\mathbb{N}^{\infty}$ to the diagram $\mathrm{U}V\colon j\mapsto\mathrm{U}V_{j}$; therefore the $v(j)$’s yield an element $v\in\mathrm{U}\lim V(S\times\mathbb{N}^{\mathrm{\infty}})$.
For $x\in\mathbb{N}\cup\{\infty\}$, the inclusion of profinite sets $S\times\{x\}\to S\times\mathbb{N}^{\infty}$ induces a morphism
$\mathrm{ev}_{x}\colon\mathrm{U}\lim V(S\times\mathbb{N}^{\mathrm{\infty}})\longrightarrow\mathrm{U}\lim V(S\times\{x\})$
and let $\mathrm{ev}_{x}^{W}$ denote the analog morphism for $\mathrm{U}W$. By construction of $v$, we have equivalences $\mathrm{ev}_{i}(v)\simeq v_{i}^{\prime}$ for $i\in I$ and $\mathrm{ev}_{\infty}(v)\simeq 0$. Using the naturality of $\mathrm{U}f\colon\mathrm{U}\lim V\to\mathrm{U}W$, we obtain equivalences
$\mathrm{ev}_{\infty}(\mathrm{U}f(v))\simeq\mathrm{U}f(\mathrm{ev}^{W}_{\infty})(v)\simeq 0$ (3.7)
in $(\mathrm{U}W)(S\times\{\infty\})$. Since $W$ is a constant sheaf, so is $\mathrm{U}W$ (using that the product in $\mathrm{U}$ is finite), hence using the description of constant sheaves provided by lemma 3.11 we see that $\mathrm{U}f(v)$ can be represented as a partition $S\times\mathbb{N}^{\infty}\cong T_{1}\amalg\dots\amalg T_{n}$ labeled by elements in $\mathrm{U}W(*)$. By the equivalence (3.7), this partition can be chosen so that $T_{n}$ contains $S\times\{\infty\}$ and is labeled by $0\in W(*)$. Since $S$ is compact, the tube lemma implies that $T_{n}$ contains an open set of the form $S\times\llbracket k,\infty\rrbracket$ for some $k\in\mathbb{N}$. Therefore we obtain equivalences
$\mathrm{U}f(v_{k}^{\prime})\simeq\mathrm{U}f(\mathrm{ev}_{k}(v))\simeq\mathrm{ev}_{k}(\mathrm{U}f(v))\simeq\mathrm{ev}_{k}(0)\simeq 0,$
which contradicts (3.6); this completes the proof. ∎
Proof of Proposition 3.4.
It follows directly from lemmas 3.6 and 3.8. ∎
This concludes the proof of Theorem 3.2.
@@ 4. Counter-example to fully faithfulness in the unbounded case
We now define the counterexample. Let $I$ is the filtered poset $(0\leftarrow 1\leftarrow\dots)$ and consider the pro-object $V\colon I\to\mathscr{D}(\mathsf{Cond}_{R})$ given by $V_{j}=\prod_{0\leq n\leq j}R[n]$, where the transition maps are the obvious projections.

### v2

@@ 3. $\infty$-categorical embedding in the bounded case
A different proof of Lemma 3.3 can be extracted from Section 2.2 of the recent paper [BL24]. Although the authors work in the particular case $R=\mathbb{Z}$, one can generalize their arguments to the case where $R$ is of finite type over $\mathbb{Z}$, with only minor changes. Let us sketch this alternative proof for completeness.
As above, we may assume that $I\cong\omega^{\mathrm{op}}$. The first step consists in writing $\lim V_{i}$ as an equalizer of two endomorphisms of the product $\prod_{i}V_{i}$ to reduce to proving that the morphism of $\infty$-groupoids
$\operatorname*{colim}_{i}\operatorname{Map}_{\mathscr{D}(\mathsf{Solid}_{R})}\mathopen{}\mathclose{{}\left(\bigoplus_{0\leqslant j\leqslant i}V_{j},W}\right)\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Solid}_{R})}\mathopen{}\mathclose{{}\left(\prod_{i}V_{i},W}\right)$ (3.5)
is an equivalence. By dévissage (e.g. using inductively [AG14, Proposition 2.13]), since arbitrary products and filtered colimits are exact in $\mathsf{Solid}_{R}$, we may reduce to the case where $V_{i}$ and $W$ are concentrated in a single degree, with $V_{i}$ a finitely generated projective $R$-module. Using that $R$ is of finite type over $\mathbb{Z}$ and arguing as in the proof of Lemma 3.3, we deduce that $\prod_{i}V_{i}$ is projective, therefore both mapping spaces in (3.5) are discrete. It remains to prove that the map
$\bigoplus_{i}\operatorname{Hom}_{\mathsf{Solid}_{R}}(V_{j},W)\longrightarrow\operatorname{Hom}_{\mathsf{Solid}_{R}}\mathopen{}\mathclose{{}\left(\prod_{i}V_{i},W}\right)$
is a bijection; this follows from fully faithfulness of the 1-categorical realization functor $\mathsf{Tate}_{{\aleph_{0}},R}\to\mathsf{Solid}_{R}$ (Proposition 1.5).
Remark 3.5.
The assumption that $R$ is of finite type over $\mathbb{Z}$ cannot be removed in the statement of Theorem 3.2, see [Sch22] for an explicit counterexample. We thank P. Scholze for pointing it out to us.
@@ 4. Counter-example to fully faithfulness in the unbounded case
We now define the counterexample. Let $I$ be the filtered poset $(0\leftarrow 1\leftarrow\dots)$ and consider the pro-object $V\colon I\to\mathscr{D}(\mathsf{Cond}_{R})$ given by $V_{j}=\prod_{0\leq n\leq j}R[n]$, where the transition maps are the obvious projections.
@@
Remark 4.4.
The counterexample we described above appears independently in [BL24, Section 2.2], where the authors attribute it to L. Mann.